Parth (parthkohli):

Which of them is correct? \(\Large \color {purple}{\rightarrow {a/b \over c} = {ac \over b}}\) or \(\Large \color {purple}{\rightarrow {a/b \over c} = {a \over bc} }\)

OpenStudy (anonymous):

(a/b)/c=a/(b*c)=(a/c)/b

Parth (parthkohli):

How can (a/b)/c equal (a/c)/b?

Parth (parthkohli):

@lgbasallote @UnkleRhaukus

Parth (parthkohli):

@Callisto @Mertsj

OpenStudy (mertsj):

Second one.

OpenStudy (callisto):

The second one...

OpenStudy (anonymous):

(20/2)/5=(20/5)/2

OpenStudy (lgbasallote):

they pretty much pwned it already

OpenStudy (mertsj):